# Imitation Learning for Combinatorial Optimisation under Uncertainty

### v1

@@ 7.2.2 Learning without interaction (Isolated Learning)
Full-Information ($\pi^{*o}$) 2031.34 3.63 22.19 25.77 0.5
Deterministic ($\pi^{*p}$) 1975.72 3.25 22.33 25.50 3.25
Table 5: Average performance metrics for Full-Information and Deterministic experts
Two-Stage Stochastic Expert
To employ the two-stage stochastic expert, with $F_{e}(\cdot)$ set to $\pi^{t}$, it is necessary to specify the number of future scenarios generated at each decision epoch. In this study, we use a fixed number of scenarios across all epochs. Alternative designs could allow the number of scenarios to vary over time—for example, by generating more scenarios in later stages, where the two-stage problem may be easier to solve and decisions are likely to be more critical with respect to the utilisation of remaining capacity.
We evaluate three configurations using 5, 15, and 30 scenarios. Figure˜5 illustrates the effect of the number of simulated episodes on the quality of the learned policy. The results indicate substantial performance improvements as the size of the training dataset increases. Performance gains are expected to plateau once the number of episodes exceeds approximately 2,000.
Table˜6 reports the results, showing that learning from a two-stage stochastic expert with multiple scenarios yields substantially better performance than learning from a full-information expert based on a single scenario. In particular, the learned policies significantly reduce the number of rejected patients. Among the configurations considered, the model trained with 30 scenarios performs best and achieves performance closest to that of the baseline policy $\pi^{t}$, as expected. These gains, however, come at the expense of increased training time, since multiple scenarios must be generated and a multi-scenario optimisation problem $(\textbf{{MIP}}_{\textbf{SPPA}})$ must be solved.
We further report detailed statistics on optimality gaps and solution times under the imposed 30-second time limit for expert calls in Table˜7. The table summarises results across approximately 400,000 expert calls. The majority of instances are solved to optimality or near optimality, as reflected by the 99% values. As the number of scenarios increases, a modest rise in the average optimality gap is observed. In cases where the optimality gap is non-zero, it is typically because the expert has identified the optimal solution but requires additional time to certify its optimality. Overall, the observed gaps suggest that the stochastic formulation remains relatively tractable and is capable of providing high-quality expert demonstrations.
With respect to solution times, the average solve time increases with the number of scenarios. Consequently, deploying $\pi^{t}$ with 30 scenarios for real-time decision-making would require approximately 5.09 seconds per decision on average. By contrast, the learned policy produces decisions within milliseconds. Solution times also vary across the planning horizon. Figure˜6 reports the average optimality gap and solution time by decision epoch. Relative to the overall averages, early decision epochs exhibit larger gaps and longer solution times, reflecting the greater problem size at the start of an episode. When using $\pi^{t}$ with a 30-second time limit, solution times at the beginning of an episode can approach 20 seconds on average. This highlights the importance of a fast reliable technique even more.
Learnt Policy Scenarios Average Cost
@@
Two-Stage ($\pi^{t}$) 5 1890.35 3.36 19.89 27.43 8.65
15 1822.45 3.03 20.02 27.38 32.47
30 1801.85 2.92 20.15 27.20 65.94
Table 6: Average performance metrics for Two-Stage Stochastic Expert
5 0 0 0.66 6.41
15 0.004 0.16 2.51 30
30 0.013 0.29 5.09 30
Table 7: Statistics across expert ($\pi^{t}$) calls
Figure 6: Average Gap and Average solve time across decision epoch
Aggregated Deterministic Expert
For a fixed number of scenarios, the deterministic expert, with $F_{e}(\cdot)$ set to $\pi^{*o}$, can be used to solve each scenario instance independently, after which the resulting solutions are aggregated to produce a single target action. Instance generation follows the same procedure as for the two-stage stochastic expert. Under this strategy, aggregation effectively serves as a heuristic approximation of a two-stage stochastic problem. Relative to the base configuration of Algorithm˜1, the primary modification is that the number of expert actions returned by $F_{j}(\cdot)$ is equal to the number of scenarios considered.
Figure˜7 illustrates the impact of the number of simulated episodes on the performance of the learned policy. Performance improves as the number of episodes increases during the initial phase of training. However, beyond approximately 500 episodes, additional data provide limited benefit, and in some cases, a marginal deterioration in performance is observed.
Table˜8 reports results across all scenario configurations considered. For aggregation, we convert the optimal actions into frequency vectors and use these as training targets for the ML model. This yields a smoother, continuous target representation, as opposed to the binary action targets used with the other expert configurations. Across all scenario counts, performance in terms of total cost remains broadly similar, indicating that increasing the number of scenarios offers limited additional benefit. Compared with policies learned from the two-stage stochastic expert, the aggregated deterministic policies perform worse overall.
Learnt Policy Scenarios Average Cost
Priority 1
Rejected
Priority 2
Rejected
Undesirable
Assignments
Train Time(hours)
Deterministic ($\pi^{*p}$) 5 1944.87 3.26 21.73 25.42 15.90
15 1950.09 3.28 21.72 25.55 50.63
30 1958.12 3.30 21.81 25.60 97.18
7.2.3 Learning with interaction
Full-Information and Deterministic Experts
Figure˜8 reports the average cost across learning iterations for both expert types. The horizontal axis corresponds to the learned policy $\pi_{i}$ obtained after iteration $i$. The results show a consistent decrease in average cost as additional data are incorporated through expert interaction. Performance metrics for the best learned policies are summarised in Table˜9, with the corresponding values indicated by dotted lines in the figure.
The models learned through interaction achieve lower average costs compared with those obtained via isolated learning, as reported in Table˜5. Table˜9 also reports the percentage improvement relative to the isolated learning models. In addition, policies learned with interaction reject fewer patients overall. No significant performance difference is observed between policies learned from the full-information expert and those learned from the deterministic expert.
As shown in Figure˜8, learning performance begins to saturate after approximately 200 iterations, although the best-performing models are obtained at later iterations. We would like to point out that the best models obtained through isolated learning required 4000 training episodes. Using interaction, the learned model improves on those values within 2000 episodes. This highlights the effect of interaction.
Training times are also reported in Table˜9. Under interactive learning, the training time associated with the deterministic expert more than doubles. For the full-information expert, training time increases substantially due to the use of the vanilla DAgger decision rule. Under this rule, the states encountered during learning differ from those observed by a pure full-information expert, which necessitates solving the $\textbf{{MIP}}_{\textbf{PPA}}$ at each newly encountered state.
Figure 8: Test performance of different learned models
Expert Scenarios Best Model Cost % Improvement
Priority 1
Rejected
Priority 2
Rejected
Undesirable
Assignments
Train Time(hours)
Full-Information 1 261 1951.76 3.9 3.03 22.41 26.93 16.40
Deterministic 1 371 1943.85 1.6 3.14 22.25 25.57 14.77
Table 9: Average performance metrics for the best model using Full-Information and Deterministic Expert with interaction
Two-Stage Stochastic Expert
Figure˜9 reports the test performance of the learned policies across learning iterations. Compared with the full-information and deterministic expert settings, a stronger and more consistent decreasing trend in total cost is observed as learning progresses.
Table˜10 summarises the average performance of the best-performing policies obtained for each scenario configuration. These policies consistently outperform their isolated learning counterparts reported in Table˜6, as indicated by the dotted lines. Among all configurations, the policy trained using 30 scenarios achieves the best performance. It yields a marginal improvement over the policy learned without interaction, while requiring a significantly smaller number of expert demonstrations.
Relative to the baseline policy $\pi^{t}$, the learned policy trained with 30 scenarios achieves a comparable average cost. Moreover, it rejects fewer priority 1 patients. This improvement is accompanied by a slightly higher number of priority 2 rejections and an increased number of assignments to non-preferred physicians.
Figure 9: Test performance of different learned models
Expert Scenarios Best Model Cost % Improvement
Priority 1
Rejected
Priority 2
Rejected
Undesirable
Assignments
Train Time(hours)
Two-stage Stochastic 5 389 1844.75 2.4 2.98 20.56 27.21 23.19
15 301 1804.59 0.9 2.84 20.45 27.01 59.22
30 251 1795.93 0.3 2.81 20.41 27.13 101.83
Table 10: Average performance metrics for the best model using the Two-stage Stochastic Expert with interaction
Aggregated Deterministic Expert
Figure˜10 illustrates the performance of learned policies obtained using different numbers of aggregated scenarios, while Table˜11 reports the performance of the best-performing policies. No significant performance differences are observed across the scenario counts considered. Notably, the best-performing policy is obtained using aggregation over only five scenarios. This observation is consistent with the spread across scenario configurations shown in Figure˜10. Furthermore, the learned models perform marginally better than those obtained through isolated learning, while requiring a smaller number of expert demonstrations.
Overall, the learned policies outperform those trained using the full-information and deterministic experts. However, they do not achieve the same level of performance as policies learned from the two-stage stochastic expert. In addition, increasing the number of aggregated scenarios to 15 or 30 does not yield meaningful performance improvements relative to the isolated learning results reported in Table˜8.
Figure 10: Test performance of different learned models
Expert Scenarios Best Model Cost % Improvement
Priority 1
Rejected
Priority 2
Rejected
Undesirable
Assignments
Train Time(hours)
Aggregated Deterministic 5 366 1938.69 0.3 3.11 22.13 25.34 29.16
15 359 1941.93 0.4 3.17 21.98 25.53 64.42
30 190 1939.165 0.9 3.17 21.96 25.33 116.45
7.3 Inference
The experimental study compares different expert configurations in terms of interaction and the computational effort required to learn high-quality policies. Among all configurations, the policy learned from the two-stage stochastic expert with 30 scenarios achieves the best performance, both with and without interaction. Allowing expert interaction further improves policy quality, measured in terms of cost, when the same number of training episodes (4,000) is used. A clear learning trend is observed across iterations: when the initial dataset is empty, average costs are high in the early iterations, but decrease rapidly within the first 10 iterations. Beyond this initial phase, steady improvements are observed for all expert configurations as additional data are incorporated. Performance across consecutive iterations does not decrease monotonically, as the incremental data collected at each iteration induces only marginal updates to the weights of the multilayer perceptron.
From a computational standpoint, the full-information and deterministic experts enable the collection of a large volume of lower-quality demonstrations within relatively short computation times. For the PPA, the two-stage stochastic formulation remains tractable, allowing this expert to be used effectively by selecting an appropriate number of scenarios. However, as the number of scenarios increases, the time required to compute an expert action also increases. Imposing a strict time limit may therefore compromise the quality of the suggested action, particularly for more complex scheduling and routing problems.
The aggregated deterministic expert offers an alternative to the two-stage stochastic expert. For the PPA, however, policies learned using the aggregated deterministic expert do not outperform those learned from the two-stage stochastic expert. Moreover, aggregation requires repeated expert calls, which substantially increases the overall training time. Empirical results indicate that Algorithm˜1 with an aggregated deterministic expert, using the same number of scenarios, incurs a higher computational cost than the two-stage stochastic expert for the PPA. Ultimately, the choice between these approaches depends on the characteristics of the problem under consideration. As demonstrated in [gawas2024ildagger], the aggregated deterministic expert can outperform the two-stage stochastic expert in settings where the two-stage problem is more difficult to solve and high-quality solutions cannot be obtained within the imposed time limits. In such cases, the aggregated deterministic expert represents a viable and effective alternative.
@@ 8 Conclusion
This paper examined the role of the expert in imitation learning for sequential decision problems under uncertainty. While imitation learning has become a practical alternative to reinforcement learning in combinatorial optimisation, the notion of an expert has remained loosely defined in the literature. To address this gap, we proposed a structured taxonomy that characterises experts along key dimensions, including uncertainty modelling, optimality, interaction strategy, labelling approach, and trajectory control. This taxonomy provides a unified perspective for interpreting prior work and for guiding expert design in future studies.
Building on this framework, we introduced a DAgger-based algorithm that explicitly incorporates expert design choices and enables systematic comparison across expert configurations. An extensive experimental study on the physician–patient assignment problem illustrates how expert properties affect learning performance, data efficiency, and computational cost. The results reveal clear trade-offs between expert quality and scalability. Full-information and deterministic experts allow rapid collection of demonstrations but often yield lower-quality policies, whereas two-stage stochastic experts provide stronger supervision at the cost of increased computation.
The experiments further show that iterative interaction with the expert can improve the policy quality, particularly during early learning iterations. Although performance improvements are not strictly monotonic, the overall trend confirms the benefit of interaction in mitigating distribution shift. Among all configurations, policies learned from a two-stage stochastic expert with a sufficiently large number of scenarios achieve the strongest performance, closely matching optimisation-based baselines while enabling significantly faster decision making.
The study also evaluated an aggregated deterministic expert as a heuristic alternative to stochastic optimisation. While this approach can be useful when two-stage problems are difficult to solve within time limits, the results indicate that it does not outperform direct stochastic supervision for the considered application and leads to higher training costs due to repeated expert queries. These findings emphasise that expert selection should be guided by problem structure and computational constraints.
Overall, this work clarifies the importance of expert design in imitation learning for sequential decision making. By formalising expert characteristics and empirically analysing their impact, the paper provides both conceptual guidance and practical insights for deploying learning-based decision policies in complex operational settings. Future work may explore adaptive expert-query strategies, richer stochastic expert models, and applications to larger-scale optimisation problems.
@@
Appendix A MIP formulations
@@ A.1 Deterministic MIP formulation
A.2 Two-stage Stochastic MIP formulation
Table˜13 introduces additional notation for the two-stage formulation, where variables and parameters are augmented with a scenario index $\omega$. All other variables and sets retain the same meaning as in Table˜12. We use $\omega$ to denote a scenario. In a two-stage formulation, the first patient is common across all scenarios and is denoted by $k=1$.
@@ A.2 Two-stage Stochastic MIP formulation
Appendix B Dirichlet-based eligibility construction

### v2

@@ 6.2.2 Performance Evaluation
Two-stage Stochastic ($\hat{\pi}^{t}$) 5 381 1845.1 11.11 26.98 27.14 20.95
15 376 1811.2 10.55 26.68 27.44 23.01
30 370 1786.0 9.8 27.15 27.2 26.56
Figure˜8 further illustrates the average MIP gaps and the time required by the expert to compute target actions at each decision epoch. The plateau observed in the gap curve indicates the decision epochs for which the approximation criterion $\kappa$ becomes binding. This effect is mirrored by a noticeably flatter profile in the average solve time across all three scenario settings. As a result, the overall training time is substantially reduced.
Figure 7: Average Gap and Average solve time at each epoch with MIP gap 0.2% across scenarios
Figure 8: Test performance of learnt models for different MIP gap settings
Additionally, we varied the MIP gap parameter when training exclusively with the two-stage stochastic expert using 30 scenarios. Specifically, we considered four values: $0.01$, $0.02$, $0.05$, and $0.1$. The corresponding test performance is shown in Figure˜8. We observe similar learning behaviour for gap limits of $0.01$, $0.02$, and $0.05$, whereas the $0.1$ setting exhibits a noticeable decline in performance, suggesting that the resulting demonstrations are of lower quality. The best-performing policies under each configuration are summarised in Table˜10. These results indicate that, although the two-stage stochastic problem can be computationally demanding, effective policies can still be learned by moderately relaxing optimality. Additional experimental results exploring variations of other algorithmic parameters are provided in the appendix.
Table 10: Average performance metrics for the best model with interaction with 30 scenarios
Expert MIP Gap Best Model Cost
@@
Two-stage Stochastic ($\hat{\pi}^{t}$) 0.01 377 1783.5 9.93 26.94 27.15 33.41
0.02 370 1786.0 9.8 27.15 27.2 26.56
0.05 341 1788.2 9.43 27.82 27.11 24.97
0.1 397 1810.7 8.82 29.3 24.85 24.84
7 Open Research Directions
Despite rapid progress in imitation learning for CO under uncertainty, several important research directions remain open.
  1. 1.
    Adaptive Expert Query Strategies — Although DAgger-style interaction improves robustness by mitigating compounding errors, expert queries remain computationally expensive in many optimisation settings. An important open direction is the development of principled mechanisms that decide both when to query and which expert to consult, based on criteria such as model uncertainty, state novelty, or expected performance gain. Integrating selective sampling and confidence-aware querying [zhang2016query, menda2019ensembledagger] from classical IL literature into stochastic optimisation frameworks is particularly promising.
  2. 2.
    Multi-Expert Supervision — The taxonomy shows that experts offer complementary strengths: deterministic experts provide efficiency, whereas stochastic experts deliver stronger supervision. As illustrated in Algorithm˜3, heterogeneous experts can be combined adaptively based on the observed state. Learning meta-policies for state-dependent expert selection could reduce computational burden while maintaining solution quality. In the PPA, early decision epochs are especially demanding for two-stage stochastic models; employing a milder expert at these stages can yield substantial time savings. A similar idea appears in ambulance routing [rautenstrauss2025optimization].
  3. 3.
    Learning under Structured Expert Imperfection — Most studies assume bounded optimality gaps or heuristic approximations but offer limited theoretical insight into how expert suboptimality propagates through imitation learning. Establishing performance guarantees that relate expert approximation quality to learned policy performance remains an important open problem, particularly in stochastic settings.
  4. 4.
    Sequential Prediction Architectures — Decision spaces are large in CO problems, making accurate end-to-end prediction difficult. A promising approach is to decompose decisions into sequential stages, with separate models predicting different components. For example, in the Dynamic Inventory Routing Problem [greif2024combinatorial], one may first select customers and then determine their visiting order. Training coordinated models, each guided by a specialised expert, may improve scalability and predictive performance.
@@ 8 Conclusion
This study examined the role of the expert in imitation learning for SDPs under uncertainty. Using the PPA case study, we compared expert configurations in terms of interaction, optimality, and computational effort. Policies trained with the two-stage stochastic expert using 30 scenarios achieved the strongest performance, both with and without interaction, while interaction further improved policy quality. Deterministic and full-information experts enable rapid generation of demonstrations but tend to produce lower-quality supervision, whereas two-stage stochastic experts provide stronger guidance at higher computational cost.
For the PPA, the two-stage formulation remains tractable with a moderate number of scenarios, but computational effort increases substantially in early decision epochs. Relaxing optimality requirements via an MIP gap limit significantly reduced training time without degrading learned policy performance, highlighting the trade-off between expert quality and scalability.
From a computational perspective, learned ML policies amortise the cost of expert computation, producing fast and stable decisions at deployment. Heuristic multi-scenario optimisation methods can reduce computation but may be biased towards sampled scenarios and perform poorly under rare events. In contrast, learned policies aggregate information across instances, improving robustness and efficiency. The Aggregated Deterministic Expert, which solves scenarios independently and aggregates decisions [gawas2024ildagger, greif2024combinatorial], offers a practical compromise: while not matching two-stage stochastic performance, it consistently outperforms single-scenario deterministic and full-information experts.
Overall, the findings emphasise the central role of expert design in imitation learning for sequential decision making. By formalising expert characteristics through a structured taxonomy and embedding them within a DAgger-based framework, this work provides a unified perspective for interpreting prior results and guiding expert selection. Future work may further explore adaptive expert-query strategies, richer stochastic expert models, and extensions to larger-scale optimisation problems.
@@
Appendix
Appendix A Background
A.1 MDP
Given the core components from Section˜2, we further describe the value function and policy in detail here.
  • Value function - The value function represents the cumulative cost over the planning horizon. For a cost-minimisation objective, the optimal value function from an initial state $x_{0}$ is given by
    $V^{*}(x_{0})=\min_{a}\mathbb{E}\Bigg\{\sum_{k=0}^{K}C_{k}(x_{k},a_{k})\,\big|\,x_{0}\Bigg\},$ (12)
    We omit the discount factor for simplicity.
  • Policy ($\pi$) - A policy $\pi$ is a mapping from states to admissible actions $a_{k}=\pi(x_{k})$. Policies may be represented in tabular form or as parameterised functions, for example, with parameter $\theta$. The value of a policy $\pi$ is defined as
    $V^{\pi}(x_{0})$ $=\mathbb{E}\Bigg\{\sum_{k=0}^{K}C_{k}(x_{k},\pi(x_{k}))\,\big|\,x_{0}\Bigg\}$ (13)
    The optimal policy $\pi^{*}$ is then defined as
    $=\operatorname*{arg\,min}_{\pi}V^{\pi}(x_{0}).$ (15)
    In the MDP literature, $V^{\pi^{*}}(x_{k})$ is referred to as the optimal cost-to-go from state $x_{k}$ at epoch $k$.
Given a value function, the optimal policy can be recovered via one-step greedy minimisation. Value functions for a fixed policy can be computed using dynamic programming methods such as value iteration or policy iteration [puterman2014markov]. While these approaches yield exact solutions, they are often computationally infeasible for large-scale problems due to the curse of dimensionality [powell2007approximate].
In some applications, interest lies only in the optimal action at a specific epoch $k$. The optimal action is given by
$a_{k}^{*}=\operatorname*{arg\,min}_{a_{k}}\Big\{C_{k}(x_{k},a_{k})+\mathbb{E}[V^{*}_{k+1}(x_{k+1})]\,\big|\,x_{k}\Big\},$ (16)
where $V^{*}_{k+1}(x_{k+1})$ denotes the optimal value of the subsequent state.
Appendix B MIP formulations
@@ B.1 Deterministic MIP formulation
B.2 Two-stage Stochastic MIP formulation
Table˜12 introduces additional notation for the two-stage formulation, where variables and parameters are augmented with a scenario index $\omega$. All other variables and sets retain the same meaning as in Table˜11. We use $\omega$ to denote a scenario. In a two-stage formulation, the first patient is common across all scenarios and is denoted by $k=1$.
@@ B.2 Two-stage Stochastic MIP formulation
Appendix C Dirichlet-based eligibility construction
@@
Appendix D Results
We present some results on additional experiments.
D.1 Aggregated Deterministic Expert
The aggregated deterministic expert [gawas2024ildagger], [greif2024combinatorial] is motivated by settings in which solving a full multi-scenario stochastic optimisation problem is computationally expensive or unstable under tight time limits. Instead of jointly optimising across scenarios, this approach solves multiple deterministic instances independently—each corresponding to a sampled future scenario—and then aggregates the resulting actions. This decomposition reduces solver complexity, enables parallelisation, and provides robustness by smoothing over scenario-specific noise. Aggregated deterministic experts therefore offer a practical compromise between purely deterministic and fully stochastic experts, especially when high-quality stochastic solutions are difficult to obtain within the available computational budget.
D.1.1 Learning without interaction
For a fixed number of scenarios, the deterministic expert, with $F_{e}(\cdot)$ set to $\pi^{*f}$, can be used to solve each scenario instance independently, after which the resulting solutions are aggregated to produce a single target action. Instance generation follows the same procedure as for the two-stage stochastic expert. Under this strategy, aggregation effectively serves as a heuristic approximation of a two-stage stochastic problem. Relative to the base configuration of Algorithm˜3, the primary modification is that the number of expert actions returned by $F_{j}(\cdot)$ is equal to the number of scenarios considered.
Figure˜9 illustrates the impact of the number of simulated episodes on the performance of the learned policy. Performance improves as the number of episodes increases during the initial phase of training. However, beyond approximately 500 episodes, additional data provide limited benefit, and in some cases, a marginal deterioration in performance is observed.
Figure 9: Aggregated Deterministic Expert
Table˜13 reports results across all scenario configurations considered. For aggregation, we convert the optimal actions into frequency vectors and use these as training targets for the ML model. This yields a smoother, continuous target representation, as opposed to the binary action targets used with the other expert configurations. Across all scenario counts, performance in terms of total cost remains broadly similar, indicating that increasing the number of scenarios offers limited additional benefit. Compared with policies learned from the two-stage stochastic expert, the aggregated deterministic policies perform worse overall.
Learnt Policy Scenarios Average Cost
Priority 1
Rejected%
Priority 2
Rejected%
Undesirable
Assignments%
Train Time (hours)
Deterministic ($\pi^{*d}$) 5 1944.87 11.8 29.0 25.4 15.90
15 1950.09 11.9 29.0 25.5 50.63
30 1958.12 12.0 29.0 25.4 97.18
D.1.2 Learning with interaction
- Figure˜10 illustrates the performance of learned policies obtained using different numbers of aggregated scenarios, while Table˜14 reports the performance of the best-performing policies. No significant performance differences are observed across the scenario counts considered. Notably, the best-performing policy is obtained using aggregation over only five scenarios. This observation is consistent with the spread across scenario configurations shown in Figure˜9. Furthermore, the learned models perform marginally better than those obtained through isolated learning, while requiring a smaller number of expert demonstrations.
Figure 10: Aggregated Deterministic Expert
Overall, the learned policies marginally outperform those trained using the full-information and deterministic experts. However, they do not achieve the same level of performance as policies learned from the two-stage stochastic expert. In addition, increasing the number of aggregated scenarios to 15 or 30 does not yield meaningful performance improvements relative to the isolated learning results reported in Table˜13.
Expert Scenarios Best Model Cost % Improvement
Priority 1
Rejected
Priority 2
Rejected
Undesirable
Assignments
Train Time(hours)
Aggregated Deterministic 5 366 1938.69 0.3 11.2 29.5 25.3 29.16
15 359 1941.93 0.4 11.5 29.3 25.5 64.42
30 190 1939.165 0.9 11.4 29.3 25.3 116.45
D.1.3 Inference
Overall aggregated deterministic expert provides an alternative to the two-stage stochastic expert, but for the PPA, it does not yield superior policies. Moreover, aggregation requires repeated expert calls, significantly increasing training time. Empirical results show that, for the same number of scenarios, Algorithm˜3 with an aggregated deterministic expert incurs a higher computational cost than the two-stage stochastic expert. That said, as demonstrated in [gawas2024ildagger], aggregated deterministic experts can outperform two-stage stochastic experts in settings where the latter are difficult to solve to high quality within practical time limits.